### v2

@@ 4.4 Local density estimator in structures using Voronoi tessellations
In order to take advantage of our MUSE dataset and of its high spectroscopic redshifts completeness (see Sect. 3.3), we compute an additional local density for each galaxy in any identified structure based on a similar but simpler approach than the above VMC method (Sect. 4.3). For each group, we perform a triangulation from the position of all galaxies with a secure spectroscopic redshift. Then Voronoi cells are allotted to each galaxy. Since each cell contains only one galaxy, the corresponding galaxy is attributed a density per unit area equals to the inverse of the Voronoi cell proper area in Mpc at the redshift of the group. This method suffers however from edge effects since tiles for galaxies at the edge of the group or of the MUSE FoV have undefined boundaries. To partially overcome this, we use the COSMOS- catalog (Khostovan et al. in prep, see Sect. 3.2) outside the MUSE FoV to add constraints on those ’edge’ galaxies. We add all galaxies with a secure (confidence flag at least 2) spectroscopic redshift in the COSMOS- catalog within the total redshift range of the group, which boundaries are defined as the minimum and maximum redshifts of members identified in MUSE data. An example of such a density map is shown for CGr32 in Fig. 9. The overall density pattern displays a good qualitative match with respect to that of the VMC method in Fig. 8.
@@ 4.6 Analysis of MAGIC groups
There is a clear correlation between the richness of groups and their Virial mass estimated from dispersion (Eq. 5), as illustrated in Fig. 11. This figure also shows that groups above redshift are mostly low-mass and low-richness. It may be that their richness is underestimated due to their lowest completeness at low mass compared to groups at . We further investigate the link between the group mass, the fraction of quiescent galaxies and the density. The population of quiescent galaxies is inferred from a color-color diagram as detailed hereafter in Sect. 5.1. The density can be estimated either from the VMC method, e.g., using the largest overdensity in the group, from the Voronoi tessellations method, e.g., using the 90 percentile of the Voronoi density in the group, or , the projected galaxy number density within a circle of radius equal to the distance to the fifth closest member to the group center (similar to the definition of Wang et al., 2020). Those three density estimators provide fairly comparable results, and we show the results obtained with the VMC method in Fig. 12. This figure shows that massive groups are also those with the densest regions, though the most massive structure is not the one with the highest density. Over the 15 most massive groups (Virial mass ), 13 are the main groups we targeted, and one among the two groups that were not targeted has only five members, making the estimate of its dynamical mass less robust. Only one targeted group (CGr23, ) has a lower dynamical mass. This may be due to the fact that this field was slightly off-centered with respect to the actual center of the group owing to our initial observing strategy that aimed at optimizing the density of star-forming galaxies within one MUSE field (see Sect. 2.1). Indeed, the center determined by Knobel et al. (2012) is ″ ( kpc) away from the center derived from the MAGIC dataset. Knobel et al. (2012) further inferred a velocity dispersion of 367 km s and a dynamical mass of for this group. The velocity dispersion we measure ( km s) may thus be that of a sub-group of star-forming galaxies, and may underestimate both the radius and dynamical mass of the whole group, which could be the reason for the position of this group in Fig. 12. The fraction of quiescent galaxies is also clearly larger for groups more massive than , where the local density is the highest (see further discussion in Sect. 5.1). The distribution of groups in these diagrams is due to combination of targeting massive groups and of searching blindly other structures within the MUSE FoV. Indeed, blind searches lead to a higher probability to find structures with less than ten members than massive groups. We also find a trend (not shown here) that the mass of a group is correlated to the mass of its most massive members.
@@ 5.1 Red sequence and environment
We identify 177 galaxies in the red sequence based on the two-color criterion defined in Ilbert et al. (2013) to select the quiescent population: and . Rest-frame colors are estimated from absolute magnitudes inferred from the Cigale run (see Sect. 2.4). This method has the advantage that it minimizes the effect of extinction and that it is less prone to evolution with redshift compared to a selection based on the specific star formation rate (sSFR, estimated in yr). As a sanity check, we display in Fig. 13 the position of quiescent galaxies in diagram showing that red-sequence galaxies have, as expected, lower SFR compared to galaxies on the main sequence, and that most quiescent galaxies have their sSFR lower than yr. The fraction of quiescent galaxies is 81% among galaxies with and only 1% with . Most of these galaxies have stellar masses above , where completeness is above 70% in MAGIC (Sect. 3.3).
@@ 6.1 An overdensity spanning CGr32, CGr34, CGr35 fields?
CGr32, CGR34 and CGr35 fields are spatially close to each others and have their main identified group in the COSMOS Wall at , , and , respectively (see Fig. 1 and Tables 1 and 4). On the one hand, the angular separation between CGr32 (MGr65) and CGr34 (MGr67) and between CGr32 and CGr35 (MGr66) is only of 4.2′ ( Mpc at ) and 6.4′ ( Mpc at ) respectively, which is less than about twice the radius Mpc of CGr32. On the other hand, the scaled redshift separation between those groups is , corresponding to only km s, and , corresponding to only km s, respectively. The redshift distribution of galaxies in CGr32 includes those of both CGr34 and CGr35. In the 20k zCOSMOS group catalog of Knobel et al. (2012), some galaxies, mainly those with photometric redshifts only, have non null probabilities to be part of several of these groups, due to the method used to assign these galaxies to groups and to the large uncertainty on their actual redshift. Thanks to our MUSE data, we are able to assign a secure spectroscopic redshift to a significant fraction of these galaxies in the MAGIC fields. We find no photometric candidate of CGr32 in the CGr34 or CGr35 fields, no candidate of CGr35 in the CGr32 or CGr34 fields, and no candidate of CGr34 in the CGr35 field. But several photometric candidates of CGr34 are in the CGr32 field. All of these have a higher probability to be part of CGr32. This raises the question about CGr32, CGr34 and CGr35 being one unique structure. CGr34 and CGr35 could also be either off-centered overdensities of the CGr32 cluster or groups infalling along filaments onto the CGr32 cluster, which seems to be supported by the filamentary structure of the COSMOS Wall (Iovino et al., 2016; Darragh Ford et al., 2019). A wider field coverage with MUSE would be needed to clearly distinguish those options.
@@ 6.2 Close groups in CGr84
The CGr84 mosaic exhibits two groups, MGr55 and MGr58 (CGr84) close in redshift space (see Fig. 19 and Table 4) at and respectively and both quite rich (39 and 33 members respectively). To assess whether those two groups may be gravitationally linked, we estimate on the one hand that their velocity difference is km s based on their scaled redshift difference , and on the other hand that their proper longitudinal separation is Mpc if the redshift difference is only due to expansion. The velocity dispersions of both structures are similar and slightly lower than 500 km s, which corresponds to a Virial radius of about 1 Mpc and a halo mass of about . Assuming that the velocity difference is not due to expansion, having the two groups gravitationally bound would require their actual separation to be lower than about 120 kpc so that their velocity difference gets lower than the escape velocity. This is much smaller than both their Virial radii and their transverse extent and is thus unlikely. On the other hand, structures with such masses and proper longitudinal separation cannot be bound because the Universe expands faster than the escape speed at a distance of 29 Mpc, which is lower than 200 km s. Another alternative would be that the two groups are located within a unique dark matter halo of much larger mass and scale, but such a scenario is unlikely. These groups are therefore most likely two distinct structures along the line-of-sight.
@@ 6.3 Extended ionized gas nebulae in CGr172
We report here the finding of a kpc bridge of ionized gas seen in [O ii] between two members of the group CGr172 (MGr57) at . This group contains 22 members and has a mass of . We show in Fig. 20 the F814W HST-ACS image as well as the contours of the [O ii] doublet emission at the redshift of the group. This ionized gas nebula has a tail shape that connects galaxies 104_CGr172 on the eastern side, to 84_CGr172 on the western side, separated by a distance of 93 kpc and having a velocity difference of about 126 km s. The latter is a diffuse low-mass () and star-forming () galaxy, composed of several clumps in the HST image. It is characterized by a high activity of star formation for its stellar mass which puts it above the main sequence relation. It seems that more small clumps are embedded in the tail between the two objects. Interestingly, 104_CGr172 appears to be in pair with 99_CGr172, with a projected proper separation of 24 kpc and a velocity difference of about 154 km s. These two galaxies are both on the red sequence and are the third and fourth most massive galaxies of the group, with stellar masses of (104_CGr172) and (99_CGr172). The group center is close to 104_CGr172, both using the barycenter and the X-ray estimates.
@@ 6.4 Quasar illuminated nebula in CGr84
A quasi stellar object (QSO) is located within field CGr84-M1 at a redshift , in group CGr84 (MGr58, ), corresponding to the source 231_CGr84. This QSO was previously reported in Prescott et al. (2006) as SDSS J100012.91+023522, and in Heintz et al. (2016) as Q 100012.9+023522.8. This galaxy has a stellar mass inferred from SED fitting of , and it appears star-forming though strongly bulge-dominated. However, these properties are not reliable because the QSO source biases both the morphology and the spectrum of the host galaxy. It has a velocity of 545 km s within the group, larger than the dispersion of CGr84 ( km s), and is located at a distance of 336 kpc from the center of the group, which is about half kpc of CGr84. The probability for this galaxy to be the brightest central galaxy of the group is thus low. The QSO spectrum displays bright and broad emission lines and is much brighter in [O iii], H, or Mg ii than in [O ii]. It is surrounded by an ionized gas nebula shining in [O ii], but brighter and more extended in [O iii]5007 (see Fig. 21), which suggests that the surrounding gas is ionized by the QSO. The nebula covers almost 10″ from South to North ( kpc in proper distance at the redshift of the QSO). It mostly extends to the North-East161616A faint galaxy not identified in the COSMOS2015 photometric catalog is found from blind search at the redshift of the group at this location, probably due to the presence of the nebula. This may be a fake source., while there is a gap between the southernmost part of the nebula and the QSO. This gap may be due to 213_CGr84, a bright and massive foreground galaxy at a redshift bridging this gap, that is thus not a member of CGr84. Indeed, this galaxy may induce noise in the data and partially screen the nebula behind it. The nebula is therefore not associated to any bright galaxy except the QSO itself. The corresponding X-ray group W11 (Gozaliasl et al., 2019) may actually be partly associated to this QSO.
@@
We estimate the monochromatic luminosity at 5100 Å by fitting the H line in the QSO spectrum, following the method described in Shen et al. (2011). We further apply a bolometric correction (Richards et al., 2006) to infer a bolometric luminosity of . The H FWHM based on a joint narrow plus broad line fit including a power-law continuum and iron complex template from Boroson & Green (1992) is km s. Using the single-epoch virial estimator from Shen et al. (2011), this results in an estimated super massive black hole mass (SMBH) mass of which would mean that the QSO is accreting at a rate of .
@@ 6.5 Gravitational arc in CGr32
A gravitational arc (514_CGr32) is clearly detected in the field CGr32-M3 around the galaxy 152_CGr32 at . This galaxy is passive and is the most massive of cluster CGr32 (MGr65) with a stellar mass of . This arc was already identified by Faure et al. (2008) as COSMOS5939+3044 and is composed of two elongated images of the same galaxy North of 152_CGr32. Thanks to the unambiguous detection of the asymmetric Ly line with MUSE, we are able to determine that the arc source is at , and to find two other images of this galaxy at the same redshift (513_CGr32 and 530_CGr32) on the other side of the lens, as illustrated in Fig. 22. The large angular separation between the lensed images indicates the presence of massive structure in addition to the lens. We model the lensing configuration of this system using a single isothermal elliptical potential at the location of 152_CGr32 (parameterized by its velocity dispersion , ellipticity and angle ), together with an external shear (parameterized with an amplitude and shear angle ). We are able to recover the locations of the four lensed images with an rms of 0.04″. The velocity dispersion of the main lens is km s, and we find a significant shear and °. The value of found can be compared with the one inferred from the MUSE spectrum. We run the penalized pixel-fitting (pPXF) code171717http://www-astro.physics.ox.ac.uk/~mxc/software/ (Cappellari & Emsellem, 2004) on the lens integrated spectrum to fit its stellar continuum and infer its velocity dispersion. The spectrum is extracted over a 6″ wide square aperture. We use the MILES stellar population synthesis library (Sánchez-Blázquez et al., 2006; Vazdekis et al., 2010; Falcón-Barroso et al., 2011) and take into account the lower spectral resolution ( Å) of this stellar library, compared to MUSE, taking advantage of the intrinsically large dispersion of the galaxy, to estimate that the velocity dispersion of the lens is about 250 km s. This value shows a good agreement with the results from our lens model. The orientation of the shear hints to the presence of a very massive structure at a position angle of ° (blue wedge in Fig. 22). This direction is in fair agreement of about ° at 1- with the direction of CGr32 center found in X-ray and from MUSE spectroscopic redshifts (39° and 44° respectively, green wedge in Fig. 22). This global agreement seems to indicate that the shear is dominated by CGr32, and the small offset may be due to a nearby less massive sub-structure or to some asymmetry in CGr32 producing a local tilt of the external shear at the location of the lens. We also determine that the velocity dispersion of this massive structure is between 800 km s and 1200 km s, which is consistent with the velocity dispersion of CGr32 of km s.

### v3

@@ 4.4 Local density estimator in structures using Voronoi tessellations
In order to take advantage of our MUSE dataset and of its high spectroscopic redshifts completeness (see Sect. 3.3), we compute an additional local density for each galaxy in any identified structure based on a similar but simpler approach than the above VMC method (Sect. 4.3). For each group, we perform a triangulation from the position of all galaxies with a secure spectroscopic redshift. Then Voronoi cells are allotted to each galaxy. Since each cell contains only one galaxy, the corresponding galaxy is attributed a density per unit area equals to the inverse of the Voronoi cell proper area in Mpc2 at the redshift of the group. This method suffers however from edge effects since tiles for galaxies at the edge of the group or of the MUSE FoV have undefined boundaries. To partially overcome this, we use the COSMOS- catalog (Khostovan et al. in prep, see Sect. 3.2) outside the MUSE FoV to add constraints on those ’edge’ galaxies. We add all galaxies with a secure (confidence flag at least 2) spectroscopic redshift in the COSMOS- catalog within the total redshift range of the group, which boundaries are defined as the minimum and maximum redshifts of members identified in MUSE data. An example of such a density map is shown for CGr32 in Fig. 9. The overall density pattern displays a good qualitative match with respect to that of the VMC method in Fig. 8.
@@ 4.6 Analysis of MAGIC groups
There is a clear correlation between the richness of groups and their Virial mass estimated from dispersion (Eq. 5), as illustrated in Fig. 11. This figure also shows that groups above redshift are mostly low-mass and low-richness. It may be that their richness is underestimated due to their lowest completeness at low mass compared to groups at . We further investigate the link between the group mass, the fraction of quiescent galaxies and the density. The population of quiescent galaxies is inferred from a color-color diagram as detailed hereafter in Sect. 5.1. The density can be estimated either from the VMC method, e.g., using the largest overdensity in the group, from the Voronoi tessellations method, e.g., using the 90th percentile of the Voronoi density in the group, or , the projected galaxy number density within a circle of radius equal to the distance to the fifth closest member to the group center (similar to the definition of Wang et al., 2020). Those three density estimators provide fairly comparable results, and we show the results obtained with the VMC method in Fig. 12. This figure shows that massive groups are also those with the densest regions, though the most massive structure is not the one with the highest density. Over the 15 most massive groups (Virial mass ), 13 are the main groups we targeted, and one among the two groups that were not targeted has only five members, making the estimate of its dynamical mass less robust. Only one targeted group (CGr23, ) has a lower dynamical mass. This may be due to the fact that this field was slightly off-centered with respect to the actual center of the group owing to our initial observing strategy that aimed at optimizing the density of star-forming galaxies within one MUSE field (see Sect. 2.1). Indeed, the center determined by Knobel et al. (2012) is ″ ( kpc) away from the center derived from the MAGIC dataset. Knobel et al. (2012) further inferred a velocity dispersion of 367 km s-1 and a dynamical mass of for this group. The velocity dispersion we measure ( km s-1) may thus be that of a sub-group of star-forming galaxies, and may underestimate both the radius and dynamical mass of the whole group, which could be the reason for the position of this group in Fig. 12. The fraction of quiescent galaxies is also clearly larger for groups more massive than , where the local density is the highest (see further discussion in Sect. 5.1). The distribution of groups in these diagrams is due to combination of targeting massive groups and of searching blindly other structures within the MUSE FoV. Indeed, blind searches lead to a higher probability to find structures with less than ten members than massive groups. We also find a trend (not shown here) that the mass of a group is correlated to the mass of its most massive members.
@@ 5.1 Red sequence and environment
We identify 177 galaxies in the red sequence based on the two-color criterion defined in Ilbert et al. (2013) to select the quiescent population: and . Rest-frame colors are estimated from absolute magnitudes inferred from the Cigale run (see Sect. 2.4). This method has the advantage that it minimizes the effect of extinction and that it is less prone to evolution with redshift compared to a selection based on the specific star formation rate (sSFR, estimated in yr-1). As a sanity check, we display in Fig. 13 the position of quiescent galaxies in diagram showing that red-sequence galaxies have, as expected, lower SFR compared to galaxies on the main sequence, and that most quiescent galaxies have their sSFR lower than yr-1. The fraction of quiescent galaxies is 81% among galaxies with and only 1% with . Most of these galaxies have stellar masses above , where completeness is above 70% in MAGIC (Sect. 3.3).
@@ 6.1 An overdensity spanning CGr32, CGr34, CGr35 fields?
CGr32, CGR34 and CGr35 fields are spatially close to each others and have their main identified group in the COSMOS Wall at , , and , respectively (see Fig. 1 and Tables 1 and 4). On the one hand, the angular separation between CGr32 (MGr65) and CGr34 (MGr67) and between CGr32 and CGr35 (MGr66) is only of 4.2′ ( Mpc at ) and 6.4′ ( Mpc at ) respectively, which is less than about twice the radius Mpc of CGr32. On the other hand, the scaled redshift separation between those groups is , corresponding to only km s-1, and , corresponding to only km s-1, respectively. The redshift distribution of galaxies in CGr32 includes those of both CGr34 and CGr35. In the 20k zCOSMOS group catalog of Knobel et al. (2012), some galaxies, mainly those with photometric redshifts only, have non null probabilities to be part of several of these groups, due to the method used to assign these galaxies to groups and to the large uncertainty on their actual redshift. Thanks to our MUSE data, we are able to assign a secure spectroscopic redshift to a significant fraction of these galaxies in the MAGIC fields. We find no photometric candidate of CGr32 in the CGr34 or CGr35 fields, no candidate of CGr35 in the CGr32 or CGr34 fields, and no candidate of CGr34 in the CGr35 field. But several photometric candidates of CGr34 are in the CGr32 field. All of these have a higher probability to be part of CGr32. This raises the question about CGr32, CGr34 and CGr35 being one unique structure. CGr34 and CGr35 could also be either off-centered overdensities of the CGr32 cluster or groups infalling along filaments onto the CGr32 cluster, which seems to be supported by the filamentary structure of the COSMOS Wall (Iovino et al., 2016; Darragh Ford et al., 2019). A wider field coverage with MUSE would be needed to clearly distinguish those options.
@@ 6.2 Close groups in CGr84
The CGr84 mosaic exhibits two groups, MGr55 and MGr58 (CGr84) close in redshift space (see Fig. 19 and Table 4) at and respectively and both quite rich (39 and 33 members respectively). To assess whether those two groups may be gravitationally linked, we estimate on the one hand that their velocity difference is km s-1 based on their scaled redshift difference , and on the other hand that their proper longitudinal separation is Mpc if the redshift difference is only due to expansion. The velocity dispersions of both structures are similar and slightly lower than 500 km s-1, which corresponds to a Virial radius of about 1 Mpc and a halo mass of about . Assuming that the velocity difference is not due to expansion, having the two groups gravitationally bound would require their actual separation to be lower than about 120 kpc so that their velocity difference gets lower than the escape velocity. This is much smaller than both their Virial radii and their transverse extent and is thus unlikely. On the other hand, structures with such masses and proper longitudinal separation cannot be bound because the Universe expands faster than the escape speed at a distance of 29 Mpc, which is lower than 200 km s-1. Another alternative would be that the two groups are located within a unique dark matter halo of much larger mass and scale, but such a scenario is unlikely. These groups are therefore most likely two distinct structures along the line-of-sight.
@@ 6.3 Extended ionized gas nebulae in CGr172
We report here the finding of a kpc bridge of ionized gas seen in [O ii] between two members of the group CGr172 (MGr57) at . This group contains 22 members and has a mass of . We show in Fig. 20 the F814W HST-ACS image as well as the contours of the [O ii] doublet emission at the redshift of the group. This ionized gas nebula has a tail shape that connects galaxies 104_CGr172 on the eastern side, to 84_CGr172 on the western side, separated by a distance of 93 kpc and having a velocity difference of about 126 km s-1. The latter is a diffuse low-mass () and star-forming () galaxy, composed of several clumps in the HST image. It is characterized by a high activity of star formation for its stellar mass which puts it above the main sequence relation. It seems that more small clumps are embedded in the tail between the two objects. Interestingly, 104_CGr172 appears to be in pair with 99_CGr172, with a projected proper separation of 24 kpc and a velocity difference of about 154 km s-1. These two galaxies are both on the red sequence and are the third and fourth most massive galaxies of the group, with stellar masses of (104_CGr172) and (99_CGr172). The group center is close to 104_CGr172, both using the barycenter and the X-ray estimates.
@@ 6.4 Quasar illuminated nebula in CGr84
A quasi stellar object (QSO) is located within field CGr84-M1 at a redshift , in group CGr84 (MGr58, ), corresponding to the source 231_CGr84. This QSO was previously reported in Prescott et al. (2006) as SDSS J100012.91+023522, and in Heintz et al. (2016) as Q 100012.9+023522.8. This galaxy has a stellar mass inferred from SED fitting of , and it appears star-forming though strongly bulge-dominated. However, these properties are not reliable because the QSO source biases both the morphology and the spectrum of the host galaxy. It has a velocity of 545 km s-1 within the group, larger than the dispersion of CGr84 ( km s-1), and is located at a distance of 336 kpc from the center of the group, which is about half kpc of CGr84. The probability for this galaxy to be the brightest central galaxy of the group is thus low. The QSO spectrum displays bright and broad emission lines and is much brighter in [O iii], H, or Mg ii than in [O ii]. It is surrounded by an ionized gas nebula shining in [O ii], but brighter and more extended in [O iii]5007 (see Fig. 21), which suggests that the surrounding gas is ionized by the QSO. The nebula covers almost 10″ from South to North ( kpc in proper distance at the redshift of the QSO). It mostly extends to the North-East161616A faint galaxy not identified in the COSMOS2015 photometric catalog is found from blind search at the redshift of the group at this location, probably due to the presence of the nebula. This may be a fake source., while there is a gap between the southernmost part of the nebula and the QSO. This gap may be due to 213_CGr84, a bright and massive foreground galaxy at a redshift bridging this gap, that is thus not a member of CGr84. Indeed, this galaxy may induce noise in the data and partially screen the nebula behind it. The nebula is therefore not associated to any bright galaxy except the QSO itself. The corresponding X-ray group W11 (Gozaliasl et al., 2019) may actually be partly associated to this QSO.
@@
We estimate the monochromatic luminosity at 5100 Å by fitting the H line in the QSO spectrum, following the method described in Shen et al. (2011). We further apply a bolometric correction (Richards et al., 2006) to infer a bolometric luminosity of . The H FWHM based on a joint narrow plus broad line fit including a power-law continuum and iron complex template from Boroson & Green (1992) is km s-1. Using the single-epoch virial estimator from Shen et al. (2011), this results in an estimated super massive black hole mass (SMBH) mass of which would mean that the QSO is accreting at a rate of .
@@ 6.5 Gravitational arc in CGr32
A gravitational arc (514_CGr32) is clearly detected in the field CGr32-M3 around the galaxy 152_CGr32 at . This galaxy is passive and is the most massive of cluster CGr32 (MGr65) with a stellar mass of . This arc was already identified by Faure et al. (2008) as COSMOS5939+3044 and is composed of two elongated images of the same galaxy North of 152_CGr32. Thanks to the unambiguous detection of the asymmetric Ly line with MUSE, we are able to determine that the arc source is at , and to find two other images of this galaxy at the same redshift (513_CGr32 and 530_CGr32) on the other side of the lens, as illustrated in Fig. 22. The large angular separation between the lensed images indicates the presence of massive structure in addition to the lens. We model the lensing configuration of this system using a single isothermal elliptical potential at the location of 152_CGr32 (parameterized by its velocity dispersion , ellipticity and angle ), together with an external shear (parameterized with an amplitude and shear angle ). We are able to recover the locations of the four lensed images with an rms of 0.04″. The velocity dispersion of the main lens is km s-1, and we find a significant shear and °. The value of found can be compared with the one inferred from the MUSE spectrum. We run the penalized pixel-fitting (pPXF) code171717http://www-astro.physics.ox.ac.uk/~mxc/software/ (Cappellari & Emsellem, 2004) on the lens integrated spectrum to fit its stellar continuum and infer its velocity dispersion. The spectrum is extracted over a 6″ wide square aperture. We use the MILES stellar population synthesis library (Sánchez-Blázquez et al., 2006; Vazdekis et al., 2010; Falcón-Barroso et al., 2011) and take into account the lower spectral resolution ( Å) of this stellar library, compared to MUSE, taking advantage of the intrinsically large dispersion of the galaxy, to estimate that the velocity dispersion of the lens is about 250 km s-1. This value shows a good agreement with the results from our lens model. The orientation of the shear hints to the presence of a very massive structure at a position angle of ° (blue wedge in Fig. 22). This direction is in fair agreement of about ° at 1- with the direction of CGr32 center found in X-ray and from MUSE spectroscopic redshifts (39° and 44° respectively, green wedge in Fig. 22). This global agreement seems to indicate that the shear is dominated by CGr32, and the small offset may be due to a nearby less massive sub-structure or to some asymmetry in CGr32 producing a local tilt of the external shear at the location of the lens. We also determine that the velocity dispersion of this massive structure is between 800 km s-1 and 1200 km s-1, which is consistent with the velocity dispersion of CGr32 of km s-1.